OpenStudy (matlee):

Let A (2, 1,) B (3, 1,) C (4, 2,) and D (1, 5). Find the component form and magnitude of the vector. ---- ---- CD - AB?

OpenStudy (anonymous):

multiply verctor CD and then substract AB from it, then we will find the magnitude

OpenStudy (matlee):

i multiply 4,2 and 1,5 for CD just like dot product?

OpenStudy (anonymous):

yes

OpenStudy (matlee):

Ok

OpenStudy (matlee):

Ok i got 35 for CD

OpenStudy (matlee):

I think i did that right , but iwas not sure if i had to multiply or add them

OpenStudy (anonymous):

u dont add them,

OpenStudy (matlee):

Ok i got 4,10

OpenStudy (matlee):

4*1 and 2*5 right

OpenStudy (anonymous):

yes

OpenStudy (matlee):

Ok, i was starting to feel like i forgot already

OpenStudy (anonymous):

wait you were right i was doing the cross product dont wanna confuse you

OpenStudy (matlee):

Ok so now i do AB which would be 2*5 =10 and 1* (-1)= -1 (10,-1)

OpenStudy (matlee):

Oh the 35 one?

OpenStudy (matlee):

Thats ok

OpenStudy (anonymous):

yeah

OpenStudy (matlee):

So AB would be 0

OpenStudy (anonymous):